### v3

# Notes on Deterministic and Stochastic Approaches in Electromagnetic Information Theory

Abstract
This paper investigates the relationship between the Number of Degrees of Freedom ($N_{DoF}\,\,$) of the field in deterministic and stochastic source models within Electromagnetic Information Theory (EIT).
Our findings demonstrate a fundamental connection between these two approaches. Specifically, we show that a deterministic model and a stochastic model with a spatially incoherent and homogeneous source yield not only the same $N_{DoF}\,\,$but also identical eigenvalues and basis functions for field representation. This key equivalence not only explains the effectiveness of deterministic approaches in EIT but also corroborates the use of classical electromagnetic methods within this new discipline.
Information and Electromagnetic Theory, Degrees of Freedom, MIMO, communication, Antennas, Electromagnetic propagation, Communication systems, Mobile communication, Radio communication, Massive MIMO, Stochastic systems
1 Introduction
In the early development of information theory, two distinct and competing paradigms emerged. One was grounded in a fundamentally physical perspective, while the other was rooted in a rigorous mathematical formalism. The foremost representative of the former was Dennis Gabor, who laid the foundations of what might be termed a "physical theory of communication" [1]. In contrast, Claude Shannon pioneered the "mathematical theory of communication" [2]. Shannon’s stochastic approach quickly became the dominant framework in modern information science, thanks to its ability to model the probabilistic nature of information.
Several decades later, the initial development of Electromagnetic Information Theory (EIT) revived a comparable dichotomy. On one side stood a non-stochastic approach, based on deterministic models drawn from classical electromagnetism, on the other, a stochastic framework that embraced the probabilistic nature of information-bearing electromagnetic fields.
However, perhaps counterintuitively, despite exhibiting a clear conceptual limitation—namely, the complete neglect of the inherently stochastic nature of the problem—since now deterministic approaches have proven so remarkably effective that they are often preferred over stochastic models. This is all the more surprising given that these methods rely on tools and results developed well before the emergence of EIT and originally intended primarily for field representation and interpolation.
This paper investigates the reasons behind the effectiveness of deterministic models used in EIT.
The literature on EIT is vast and rich of important results obtained by a large community actively working both on non stochastic approaches [3, 4, 5, 6, 7, 8, 9, 10, 11] and stochastic approaches [12, 13, 14, 15].
To provide a focused discussion, we will primarily refer to a select number of references from the extensive literature on EIT.
Our discussion of the deterministic approach is grounded in the foundational papers [16, 17, 18] and in the work reported in [19], who adapted the results of the first group to the EIT paradigm. For the stochastic approach, our primary reference is the classic text by Leonard Mandel and Emil Wolf [20].
While we will adhere to the notations of these references wherever possible, some simplifications and modifications are necessary due to the interdisciplinary nature of this paper. This work draws from fields such as coherent optics, information theory, antenna theory, and radiometry, each with its own independent notational conventions. Details that require further exploration are relegated to footnotes to maintain a continuous discussion.
While the analysis in this paper is limited to scalar free-space propagation, the approach can be extended to full vector fields by using cross-spectral density tensors [20](Sect. 6.6).
Figure 1: Geometry of the problem in case of non stochastic source; $\mathbf{r}^{\prime}\in D$, where $D$ is the the source domain while the $\mathbf{r}$ lies on the observation manifold $S_{obs}$, that in this paper will be a curve or a surface.
2 Some Recalls on the Non-Stochastic Approach to EIT
As preliminary step, let us consider the ’pure spatial communication system’ model shown in Fig. 1, consisting in an harmonic electromagnetic source $D$, whose surface area is $A_{S}$, and an observation surface $S_{obs}$ [19].
With $E^{d}(\mathbf{r})$ and $J^{d}(\mathbf{r})$ we will denoted the phasor representation of the electric field and the currents density on the source. The superscript $d$ stands for deterministic. The time dependence $e^{j\bar{\omega}t}$ will be understood and dropped.
The field $E^{d}(\mathbf{r})$ observed on the domain $S_{obs}$ is related to the current distribution ${J}^{d}(\mathbf{r})$ on the source domain $D$ by the radiation operator [19]:
${E}^{d}(\mathbf{r})=\int_{D}J^{d}(\mathbf{r}^{\prime})G(\mathbf{r},\mathbf{r}^{\prime})d\mathbf{r}^{\prime}$ (1)
where the free–space Green’s function is $G(\mathbf{r},\mathbf{r}^{\prime})=-\,j\bar{\omega}\mu_{0}\frac{e^{-j\beta R}}{4\pi R},\qquad R=|\mathbf{r}-\mathbf{r}^{\prime}|,\quad\mathbf{r}\in S_{\mathrm{obs}}$, $\beta=\bar{\omega}\sqrt{\epsilon_{0}\mu_{0}}$ is the free–space wavenumber, and $\epsilon_{0}$ and $\mu_{0}$ are the permittivity and permeability of free space, respectively. The functions $E^{d}$ and $J^{d}$ are square–integrable, and $J^{d}$ is bounded.
The radiation operator, being compact, can be expanded using Hilbert-Schmidt representation [21], obtaining a proper basis for the currents and the field
$E^{d}(\mathbf{r})$ $=\sum_{n}y^{d}_{n}{u}_{k}(\mathbf{r})$ (2)
$J^{d}(\mathbf{r}^{\prime})$ $=\sum_{n}x^{d}_{n}{v}_{k}(\mathbf{r}^{\prime})$ (3)
These basis diagonalize the radiation operator, such that $y^{d}_{n}=\sigma_{h}x^{d}_{n}$. The functions ${u}_{n}(\mathbf{r}^{\prime})$ and ${v}_{n}(\mathbf{r}^{\prime})$ are respectively the left and right singular functions and $\sigma_{n}$ are the singular values of the operator [21]. Singular values are real non negative quantities. We suppose that they are sorted in descending way, i.e. $\sigma_{1}\geq\sigma_{2}\geq$.... The number of significant singular values fixes the Number of Degrees of Fredom ($N_{DoF}\,\,$) of the field [16]. The singular functions $v_{n}(\mathbf{r})$ and $u_{n}(\mathbf{r})$ are obtained as solution of eigenvalue problems [22]. In particular, $u_{n}(\mathbf{r})$ are the eigenfunctions of the eigenproblem whose kernel is:
$K_{u}(\mathbf{r_{1}},\mathbf{r_{2}})=\int_{D}G(\mathbf{r_{1}},\mathbf{r}^{\prime})G^{*}(\mathbf{r_{2}},\mathbf{r}^{\prime})d\mathbf{r}^{\prime}$ (4)
where $G^{*}(\mathbf{r}_{2},\mathbf{r}^{\prime})$ is the adjoint of the Green’s function, which, in the case of free space and scalar Green’s function, reduces to its complex conjugate. The $n$-th eigenvalue of this eigenproblem is equal to the square of the $n$-th singular value of the Hilbert-Schmidt expansion of the radiation operator.
This approach, developed from studies on electromagnetic field representation [16, 17], makes the extensive literature on field interpolation directly applicable to EIT  [^1] For mathematical rigor, some care must be taken when moving from approximation theory to information theory. The former essentially involves a covering problem for the set of fields, while the latter requires solving a packing problem [5]. This results in a different $N_{DoF}\,\,$. However, this difference is therefore generally negligible for large $N_{DoF}\,\,$..
3 EIT model with stochastic fields
3.1 Information carried by stochastic fields
Now, with reference to the same model shown in Fig. 1, let us consider a statistical ensemble of current distributions $\{j^{(r)}(\mathbf{r}^{\prime},t)\}$, where each realization $j^{(r)}$ represents the current density at a point specified by a position vector $\mathbf{r}^{\prime}\in D$ at time $t$. This ensemble is modeled as a stationary stochastic process, at least in the wide sense.
Analogously, we consider a statistical ensemble of radiated fields $\{e^{(r)}(\mathbf{r},t)\}$, stationary, at least in the wide sense, whose single realization is the electric field at a point specified by a position vector $\mathbf{r}\in S_{obs}$ at time $t$.
In the following, we will use some notation simplifications. For quantities involving the field, we do not use a subscript, while for quantities involving the current $J$, we will use the $J$ subscript.
The cross-correlation function (also called mutual coherence) of the field at the observation points $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ with time delay $\tau$, and the cross-correlation function of the current density at the source points $\mathbf{r}^{\prime}_{1}$ and $\mathbf{r}^{\prime}_{2}$, are [20, Sec. 4.2]:
$\Gamma(\mathbf{r}_{1},\mathbf{r}_{2};\tau)$ $=\mathbb{E}\!\left\{\hat{E}(\mathbf{r}_{1},t)\,\hat{E}^{*}(\mathbf{r}_{2},t+\tau)\right\},$ (5)
$\Gamma_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2};\tau)$ $=\mathbb{E}\!\left\{\hat{J}(\mathbf{r}^{\prime}_{1},t)\,\hat{J}^{*}(\mathbf{r}^{\prime}_{2},t+\tau)\right\}.$ (6)
where $\hat{E}$ is the complex analytic representation of $e^{r}(t,\mathbf{r})$ whose imaginary part is related to the original waveform by Hilbert transformation, $\mathbb{E}\{\hat{E}\}$ denotes the ensemble average of $\hat{E}$, and $\hat{E}^{*}$ is the complex conjugate of $\hat{E}$ [20](Sect. 3). A similar notation is used for $\hat{J}$.
The discussion on the stochastic meaning of the field is analogous to that of the currents. For this reason, we will focus mainly on the field.
In the ensemble $\{\hat{E}(\mathbf{r},t)\}$, each member is a possible realization of the electric field at every point in space and at every instant in time. This ensemble accounts for all temporal fluctuations and spatial variations of the field.
Under the wide-sense stationarity (WSS) assumption for both the field and the current, the temporal correlation functions depend only on the time lag $\tau$.
The cross-spectral density (CSD) is the Fourier transform in $\tau$ [20, Sec. 4.3]:
$W(\mathbf{r}_{1},\mathbf{r}_{2},\omega)$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\Gamma(\mathbf{r}_{1},\mathbf{r}_{2};\tau)\,e^{+j\omega\tau}\,d\tau,$ (7)
$W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2},\omega)$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\Gamma_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2};\tau)\,e^{+j\omega\tau}\,d\tau.$ (8)
In the following, whenever we analyze the problem at a specific angular frequency $\omega=\bar{\omega}$, the dependence on $\omega$ will be understood and dropped, paralleling the harmonic signal representation. Accordingly, we will call $W(\mathbf{r}_{1},\mathbf{r}_{2})=W(\mathbf{r}_{1},\mathbf{r}_{2},\bar{\omega})$ and $W_{J}(\mathbf{r^{\prime}}_{1},\mathbf{r^{\prime}}_{2})=W_{J}(\mathbf{r^{\prime}}_{1},\mathbf{r^{\prime}}_{2},\bar{\omega})$
Applying the Karhunen-Loëve expansion [20](Sect. 4.3), the solution of the eigenvalue problem:
$\int_{S_{obs}}W(\mathbf{r}_{1},\mathbf{r}_{2})\psi_{n}(\mathbf{r}_{1})d\mathbf{r}_{1}$ $=\lambda_{n}\psi_{n}(\mathbf{r}_{2})$ (9)
$\int_{D}W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})\psi_{J_{n}}(\mathbf{r}^{\prime}_{1})d\mathbf{r}^{\prime}_{1}$ $=\lambda_{J_{n}}\psi_{Jn}(\mathbf{r}^{\prime}_{2})$ (10)
gives the ensemble of functions:
$E(\mathbf{r})$ $=\sum_{n}a_{n}\psi_{n}(\mathbf{r})$ (11)
$J(\mathbf{r}^{\prime})$ $=\sum_{n}a_{Jn}\psi_{Jn}(\mathbf{r}^{\prime})$ (12)
wherein $\mathbb{E}\{a_{n}(\omega)a_{m}^{*}(\omega)\}=\lambda_{n}\delta_{nm}$ and $\{\delta_{nm}\}$ is Kronecker symbol, equal to 1 if $m=n$, and 0 otherwise. The same property is valid for the coefficients $a_{J_{n}}$.
This representation has a number of interesting characteristics. The CSD $W(\mathbf{r}_{1},\mathbf{r}_{2})$ is equal to:
$W(\mathbf{r}_{1},\mathbf{r}_{2})=\mathbb{E}\{E(\mathbf{r}_{1})E(\mathbf{r}_{2})^{*}\}$ (13)
where the average is taken over an ensemble of space-frequency realizations. Accordingly, this formula is the equivalent in space domain-frequency domain of the cross correlation function in space-time domain in Eq. (5) (where the average is taken over an ensemble of space-time realizations).
The functions $\psi_{n}(\mathbf{r})$ and $E(\mathbf{r})$ verify the Helmholtz equation. This is also true for the CSD, that satisfies the two Helmholtz equations [20](Sect. 4.4):
$\nabla^{2}_{1}W(\mathbf{r}_{1},\mathbf{r}_{2})+\beta^{2}W(\mathbf{r}_{1},\mathbf{r}_{2})=0$ (14)
$\nabla^{2}_{2}W(\mathbf{r}_{1},\mathbf{r}_{2})+\beta^{2}W(\mathbf{r}_{1},\mathbf{r}_{2})=0$ (15)
where $\nabla^{2}_{1}$ and $\nabla^{2}_{2}$ operate respectively on $\mathbf{r}_{1}$ and on $\mathbf{r}_{2}$ keeping $\mathbf{r}_{2}$ and $\mathbf{r}_{1}$ fixed. Loosely speaking, cross-spectral density ’propagates’ in the space according to laws fixed by Maxwell’s equation.
The eigenfunctions $\psi_{n}(\mathbf{r})$ represent completely coherent modes. In general, $E(\mathbf{r})$ is only partially coherent. From its CSD we introduce the spectral degree of coherence [20, Sec. 4.3]:
$\mu(\mathbf{r}_{1},\mathbf{r}_{2})\;=\;\frac{W(\mathbf{r}_{1},\mathbf{r}_{2})}{\sqrt{\,W(\mathbf{r}_{1},\mathbf{r}_{1})\,W(\mathbf{r}_{2},\mathbf{r}_{2})\,}}$, (16)
where $W(\mathbf{r}_{1},\mathbf{r}_{1})$ (and similarly $W(\mathbf{r}_{2},\mathbf{r}_{2})$) is the spectral density of the electric field at the point $\mathbf{r}_{1}$, denoted $S(\mathbf{r}_{1})$ (and $S(\mathbf{r}_{2})$).
Determining mutual information generally requires knowledge of the full joint probability distribution of the variables, which implies knowing the complete statistics of the underlying process. However, in the case of zero-mean stationary complex Gaussian circularly symmetric processes, [^2] The assumption of a Gaussian distribution is widely accepted in wireless communication systems, as it maximizes Shannon entropy and information content of the field. A circularly symmetric complex Gaussian process is characterized by a probability distribution that is invariant under phase rotations. This assumption implies that the pseudo-covariance matrix (which describes correlations between non-conjugate complex variables) is identically null, thereby simplifying the process’s characterization significantly. If the process has a non-zero mean, the CSD used for mutual information calculation must be evaluated on the fluctuating part of the stochastic process. That is, instead of $E(\mathbf{r}_{1})$ and $E(\mathbf{r}_{2})$, the CSD is evaluated on $E^{\prime}(\mathbf{r}_{1})=E(\mathbf{r}_{1})-\mathbb{E}\{E(\mathbf{r}_{1})\}$ and $E^{\prime}(\mathbf{r}_{2})=E(\mathbf{r}_{2})-\mathbb{E}\{E(\mathbf{r}_{2})\}$ [23]. the statistical properties are entirely described by covariance (or correlation) function, and the differential mutual information $I_{MI}(\mathbf{r}_{1};\mathbf{r}_{2})$ between the two points is given by [24]:
$I_{MI}(\mathbf{r}_{1};\mathbf{r}_{2})=\log\left(\frac{1}{1-|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|^{2}}\right)$ (17)
This formula quantifies how much information one random variable (the field at $\mathbf{r}_{1}$) can be inferred by another (the field at $\mathbf{r}_{2}$)  [^3] The differential mutual information tends to infinity as $\mathbf{r}_{1}\rightarrow\mathbf{r}_{2}$. At $\mathbf{r}_{1}=\mathbf{r}_{2}$, the formula is undefined, and the mutual information reduces to the differential entropy of the variable itself [24]. . The lower the mutual information, the more information can be independently encoded at the two points in space.
Eqs. (16) and (17) allow to obtain the mutual information from the knowledge of CSD. In the following we will focus our attention on the estimation of CSD.
$W(\mathbf{r}_{1},\mathbf{r}_{2})=$
$\int_{D}\int_{D}W_{J}(\bm{\mathbf{r}^{\prime}}_{1},\bm{\mathbf{r}^{\prime}}_{2})G(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})G^{*}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})d\mathbf{r}^{\prime}_{1}d\mathbf{r}^{\prime}_{2}$ (18)
This equation represents the most general law for the propagation of second-order coherence.
In the following analysis, we will assume a spatially uncorrelated source  [^4] Rigorously speaking, a source cannot be perfectly spatially uncorrelated [25]. However, the correlation is subject to the low-pass filtering process of the free-space Green’s function that tends to filter the higher spatial frequencies [20] (Sect. 5.3). This implies that the approximation of a spatially incoherent source does not significantly affect the field results when the source’s intrinsic spatial correlation length is sufficiently short (let us say, shorter than $\lambda/2$), as propagation naturally increases the coherence distance. , which can be considered the ’richest source’ in terms of spatial information. For such a source, the cross-spectral density of the current is given by $W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=S_{J}(\mathbf{r}^{\prime}_{1})\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2})$, where $S_{J}(\mathbf{r}^{\prime}_{1})=W_{J}(\mathbf{r}^{\prime},\mathbf{r}^{\prime})$ is the source spectral density per unit angular frequency..
Under this hypothesis, the formulas (18) simplify as
$W(\mathbf{r}_{1},\mathbf{r}_{2})=\int_{D}S_{J}(\mathbf{r}^{\prime})G(\mathbf{r}_{1},\mathbf{r}^{\prime})G^{*}(\mathbf{r}_{2},\mathbf{r}^{\prime})d\mathbf{r}^{\prime}$ (19)
This integral is of the same form as is encountered in the evaluation of the field diffracted by an aperture $D$ in an opaque screen at a point $\mathbf{r}_{1}$ illuminated by a harmonic spherical wave centered in $\mathbf{r}_{2}$, or also of the scattered field in Born approximation. This similarity is at basis af many properties that the CSD shares with the theory of diffraction and scattering integrals.
In particular, the Van Cittert-Zernike theorem [20] (Sects. 4.4.5 and 5.5.1) states that under the Fraunhofer approximation, the CSD of a spatially incoherent planar source is proportional to the Fourier transform of its intensity distribution. This result resembles the planar near-field to far-field transformation in antenna theory, and indeed shares many common characteristics. Consequently, the Van Cittert-Zernike theorem serves as a powerful tool to connect information and electromagnetism within the Fraunhofer zone.
The $N_{DoF}\,\,$of a stochastic process can be obtained by the Karhunen-Loève expansion. With reference to Eq. (9), we recall that the eigenvalues $\lambda_{n}$ are real and non-negative, and exhibit a step-like behavior: a finite number of them are significant (close to a maximum value), while the rest are negligible. The number of these significant modes is the $N_{DoF}\,\,$of the stochastic field. It can be easily shown that the $N_{DoF}\,\,$of the field radiated by an uncorrelated source coincides with the $N_{DoF}\,\,$of the non statistical model (1) provided that we consider a "weighted" radiation operator whose kernel is $G^{\prime}(\mathbf{r},\mathbf{r}^{\prime})=G(\mathbf{r},\mathbf{r}^{\prime})\sqrt{S_{J}(\mathbf{r}^{\prime})}$.
Let us consider the case an uncorrelated homogeneous  [^5] A homogeneous source is a source whose statistical properties are uniform across its surface, i.e. whose spectral density is constant on the source. source with $S_{J}(\mathbf{r}^{\prime}_{1})$=1. In this case the eigenproblem in Eq. (9) is the same eigenproblem that gives the left singular functions of the radiation operator, whose kernel is Eq. (4). Consequently, the eigenfunctions of the CSD operator coincide with the left singular functions of the radiation operator, and the eigenvalues of the CSD operator coincide with the square of the singular values of the radiation operator, giving a direct connection in terms of $N_{DoF}\,\,$between the stochastic and non-stochastic model.
This key equivalence has a rather general validity (see Appendix A for a proof in the general case) and explains the remarkable success of classical electromagnetic methods in the EIT analysis of spatial information.
3.2 $N_{DoF}\,\,$in stochastic field
Estimation of $N_{DoF}\,\,$of stochastic fields from the CSD can be obtained by the Wigner transform[20](Sect. 5.7.2), originally introduced in quantum theory to obtain a phase-space [^6] Put simply, phase space is a multidimensional space in which each point represents a complete state of a system. quasi-probability distribution [^7] The WDF is a quasi-probability distribution since it can also assume negative values, which are related to interference phenomena. representation between the two conjugate variables, position (space) and momentum [20](Sect. 11.8.1). In the context of optics, it provides a phase-space representation between the conjugate variables of spatial position and wavenumber vector [26].
Wigner transformation is formally defined as [27]:
$\mathcal{W}(\mathbf{r},\mathbf{k})=\int W\left(\mathbf{r}+\frac{\Delta\mathbf{r}}{2},\mathbf{r}-\frac{\Delta\mathbf{r}}{2},\right)e^{-j\mathbf{k}\cdot\Delta\mathbf{r}}d(\Delta\mathbf{r})$ (20)
where $\mathcal{W}(\mathbf{r},\mathbf{k})$ is the Wigner Distribution Function (WDF), $\mathbf{r}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$ serves as the midpoint coordinate (local position), $\Delta\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ is the difference coordinate (spatial separation), and $\mathbf{k}$ is the wave vector. The CSD can be recovered from the WDF via the inverse Fourier transform.
In essence, for each spatial point $\mathbf{r}$, the Wigner function $\mathcal{W}(\mathbf{r},\mathbf{k})$ provides a "spectrum" of wavenumbers $\mathbf{k}$ that characterize the field’s coherence properties at that location. The marginal of $\mathcal{W}$ with respect to $\mathbf{k}$ yields the spatial distribution of the spectral intensity, while, the marginal with respect to position $\mathbf{r}$ yields the angular spectrum. Finally, in a predominantly radiative field configuration, integrating the WDF over the propagating part of phase space (both $\mathbf{r}$ and $\mathbf{k}$) yields a quantity proportional to the radiative spectral power [^8] Under a local plane-wave (i.e., far-field) assumption, an electric-field–based Wigner description is sufficient to estimate power density. In that case the proportionality factor is $1/\!\big(2\,\eta_{0}\,(2\pi)^{n}\big)$, where $\eta_{0}$ is the free space wave impedance , $n$ is the dimensionality of the spatial domain $\mathbf{r}$, and $(2\pi)^{n}$ follows from the Fourier-transform normalization. This proportionality does not generally hold in more complex field configurations (e.g., in the reactive near field)..
Since the Wigner transform relates conjugate variables, the localization of these variables in phase space is limited by the uncertainty principle. This fundamental limit defines a minimum ’uncertainty volume’ (called ’mode’ or ’cell’ in quantum theory) in phase space, that turns to be [^9] For a single pair of conjugate variables ($n=1$), such as position $x$ and wave vector component $k_{x}$, the uncertainty principle implies $\Delta x\Delta k_{x}\geq 1/2$. While this lower bound can be achieved using Gaussian functions [1], the Balian-Low theorem [28] states that it is not possible to construct a complete orthonormal basis for $L^{2}(\mathbb{R})$ using Gabor atoms (i.e., Gaussian basis functions) where all elements simultaneously achieve this minimum Heisenberg uncertainty. To define a complete orthonormal basis for functions in $L^{2}(\mathbb{R})$, each basis function must occupy a minimum area of $2\pi$ in the phase space, which represents the minimum ”cell” of phase space [29]. equal to to $(2\pi)^{n}$. An estimation of the $N_{DoF}\,\,$can then be obtained by counting the number of these minimum ’uncertainty volumes’ that are packed into the available phase space [^10] It is undertood that the $N_{DoF}\,\,$doubles when polarization is considered.:
$N_{DoF}\approx\frac{\text{Volume}(\text{phase space})}{(2\pi)^{n}}$ (21)
where with Volume we mean the volume of Wigner function support in phase space, and $n$ represents the number of spatial dimensions being considered for the field’s transverse extent (i.e., the dimensionality of the spatial domain $\mathbf{r}$)  [^11] For example, $n=1$ in case of 1D spatial problem (e.g., a line source), $n=2$ if considering a 2D transverse plane (e.g., a planar source or observation screen). Since for each such spatial dimension, there is a corresponding conjugate wave vector dimension, the phase space itself is $2n$-dimensional. . For example, in case of a line source $\ell$ long the $N_{DoF}\,\,$in the $z>0$ half-plane ($n=1$) is given by the the volume of the phase space, $2\beta\times\ell$, divided by $2\pi$  [^12] Note the similarity of this approach to the Landau theorem in Information Theory [30] to the identification of $N_{DoF}\,\,$. As an example, in case of a band-limited complex signal having total (bilateral) angular bandwidth $\hat{\Omega}_{bil}$ acquired over an interval $T$, the $N_{DoF}\,\,$, derived from the Landau theorem, is given by the time-bandwidth product $\hat{\Omega}_{bil}\times T$ divided by $2\pi$..
In case $n=2$, the extension of the wavenumber domain is restricted to the "visible range" [31], i.e. a circle in the wavenumber $k_{x}-k_{y}$ plane having radius $\beta$  [^13] This conditions assures that $k_{z}\in\mathbb{R}$, so that the corresponding $\mathbf{k}$ is not associated with an evanescent wave [32].. As an example, for a spherical source having surface $A_{S}$ and an observation surface placed at large distance from the source and completely surrounding it, we obtain:
$N_{DoF}\approx\frac{\pi A_{S}}{\lambda^{2}}$ (22)
These results coincides with the results obtained in EIT using non-stochastcic approaches based on modal expansion [33], the Weyl’s law [34], the classic near field-far field transformation [32], [14], and with the results based on field sampling theory [17], apart from a $\pi/4$ factor [^14] This discrepancy arises from the common sampling strategy in current electromagnetic field measurement systems, where sampling positions are constrained to lie along the coordinate curves of the scanning surface due to mechanical limitations [35]. [18].
The covering of phase space can also be translated into the space-angular domain, which roughly corresponds to the aperture-solid angle product used in antenna theory. This product represents the ’effective cross-sectional area’ of the power transfer channel through which energy flows from the source to the observation domain, accounting for both spatial extent and angular spread.
This method becomes an exact method to evaluate the $N_{DoF}\,\,$in the high-frequency asymptotic regime ($\lambda\rightarrow 0$), i.e., the geometrical optics approximation. In this regime, the $N_{DoF}\,\,$is strictly related to the étendue ($G_{e}$) [36], a pure geometrical quantity of great importance in radiometry [^15] Other interesting connections between radiometry and the stochastic approach used in this paper, that are not discussed here due to space limitations, are reported in [37]. [20](Sect. 5.7) since it is conserved under the geometrical optics approximation in the absence of diffraction, absorption, or scattering phenomena [^16] This invariance corresponds to the conservation of entropy in the absence of absorption, emission, or scattering processes [38]..
The étendue is defined as the integral over the source area $S$ and the solid angle of emission $\Omega$ of the projected area element multiplied by the differential solid angle element:
$G_{e}=\int_{S}\int_{\Omega}\cos\theta\,d\Omega\,dS$ (23)
This can be expressed as $G_{e}=A_{S}\Omega^{\prime}$, where $A_{S}$ is the source area and $\Omega^{\prime}$ is the projected solid angle, which is defined as:
$\Omega^{\prime}=\int_{\Omega}\cos\theta\,d\Omega$ (24)
and accounts for the angular distribution of the emitted radiation.
The projected solid angle, $\Omega^{\prime}$, is related to the spatial wavenumber domain by the integral:
$\Omega^{\prime}=\frac{1}{\beta^{2}}\iint_{\Xi}dk_{x}\,dk_{y}$
where $\beta=2\pi/\lambda$ is the wavenumber and $\Xi$ is the domain on the $k_{x}-k_{y}$ plane defined by the solid angle of emission. Since the maximum extent of the $k_{x}-k_{y}$ domain in the visible region is a disk of radius $\beta$, the maximum possible projected solid angle is limited to $\Omega^{\prime}\leq\pi$.
The phase-space volume must be divided by the unit cell. When the unit cell, which has an area of $(2\pi)^{2}$, is mapped to the projected angular domain, its effective size, representing one "angular resolution element," is given by:
$\Omega^{\prime}_{\text{cell}}=\frac{(2\pi)^{2}}{\beta^{2}}=\lambda^{2}.$ (25)
Consequently, the number of angular resolution elements (or unit angular cells) is:
$N\simeq\frac{A_{S}\Omega^{\prime}}{\lambda^{2}},$ (26)
where $A_{S}$ is the source area, $\Omega^{\prime}$ the emission solid angle, and $\lambda$ the wavelength. This expression reduces to
$N\simeq\frac{\pi A_{S}}{\lambda^{2}}$ (27)
in the case of a convex observation surface completely surrounding the source, where the projected solid angle $\Omega^{\prime}$ approaches its maximum value of $\pi$. In the case of non-convex sources, or when self-blocking (occlusion) effects occur, the expression $G_{e}=A_{S}\Omega^{\prime}$ may overestimate the effective étendue, as it implicitly assumes full visibility of the emitting surface in all relevant directions. In such cases, it is necessary to evaluate the effective projected area of the visible (non-occluded) portion of the source as seen from each direction in the observation domain. These results align with recent advances in $N_{DoF}\,\,$theory [10] that use a characteristic mode expansion to relate the $N_{DoF}\,\,$to the total shadow area of the source.
It is worth noting that the preceding discussion is based on radiometry, which studies the flux of energy (i.e., power in harmonic signals). This indicates that the $N_{DoF}\,\,$is a metric inherently linked to the physics of energy transfer. The ’dual role’ of the $N_{DoF}\,\,$in defining performance for both information and power density distribution in space confirms that the $N_{DoF}\,\,$in approximation theory and in information theory are two faces of the same coin. This property has been utilized in the development of antenna synthesis strategies with mixed information-energy goals [22, 39].
Figure 2: The red curves show the $\xi$ hyperbolas for a linear source of length $\ell=8\lambda$ (drawn as a green line). The gray rectangles are "receiving domains".
Figure 3: From a ’spatial information’ point of view, linear source sees the scene in Fig. 2 in a distorted space.
Figure 4: Real part of the Wigner distribution along the hyperbola passing through $\bar{\mathbf{r}}=(20\lambda,16\lambda)$; $s$ is the abscissa along the curve normalize to the wavelength, and $k_{s}$ the component tangent to the curve of the $\mathbf{k}$ vector; close to the source we have a large range of $\mathbf{k}$ vector pointing in different direction, giving a complicated ’power flow’ pattern; for large $s$ the component the vector $\mathbf{k}$ tangent to the curve tends to $\beta=2\pi/\lambda$ and $\mathbf{k}$ tends to be aligned to the curve.
4 Spatially incoherent homogeneous linear source
To clarify the concepts from the previous section, we consider a spatially incoherent linear source of length $\ell=2a$ placed along the $y$-axis. We will limit our analysis to the yzy-z plane. For the sake of notational simplicity, we will omit the variable $x$ in this section, as its value is always zero.
As a preliminary step, we recall some results from the framework of field interpolation, and in particular from the bandlimited field approximation. In [18], it is shown that, to minimize the number of spatial samples required to interpolate the field, these samples must be placed at the intersections between the observation curve and a subset of the constant-$\xi$ curves defined therein. In the specific linear source geometry considered, the constant-$\xi$ curves are hyperbolas with foci at the extremal points of the source, while the subset, which we will refer to as $Q_{\xi}$ in the following, consists of hyperbolas intersecting the source at a distance of $\lambda/2$ from each other [^17] The exact distance and number of curves required to interpolate the field depends on the required maximum interpolation error [17]..
Figure 2 illustrates these hyperbolas as red curves for a linear source of length $\ell=8\lambda$ (green line). The gray rectangles are "receiving domains" whose function is detailed in the subsequent discussion.
Now, let us consider the EIT stochastic approach. The CSD in Eq. (19) for a spatially uncorrelated, homogeneous line source with spectral density $S_{J}(\mathbf{r}^{\prime})=1$ on the segment $\mathbf{r}^{\prime}=(0,y^{\prime})$, $a\leq y^{\prime}\leq a$, and zero elsewhere, is
$W(\mathbf{r}_{1},\mathbf{r}_{2})\;=\;\frac{(\bar{\omega}\mu_{0})^{2}}{16\pi^{2}}\int_{-a}^{a}\frac{e^{+\,j\beta\,[\,R_{2}(y^{\prime})-R_{1}(y^{\prime})\,]}}{R_{1}(y^{\prime})\,R_{2}(y^{\prime})}\,dy^{\prime},$ (28)
where $\mathbf{r}_{n}=(z_{n},y_{n})$ with $z_{n}>0$ and
$R_{n}(y^{\prime})\;=\;\sqrt{\,z_{n}^{2}+(y_{n}-y^{\prime})^{2}\,},\qquad n=1,2.$ (29)
Before showing the numerical solution, it is useful to investigate some general characteristics considering a simple linear approximation of the phase valid in Fraunhofer approximation when the variation of $y^{\prime}$ is small compared to $z_{n}$:
$R_{n}(y^{\prime})\approx\frac{R_{n}^{+}+R_{n}^{-}}{2}+y^{\prime}\frac{R_{n}^{+}-R_{n}^{-}}{2a}$ (30)
where $R_{n}^{\pm}=\sqrt{z_{n}^{2}+(y_{n}\pm a)^{2}}$ are the distance from the source’s extreme points, obtaining
$W(\mathbf{r}_{1},\mathbf{r}_{2})$ $\;\approx\;\frac{(\bar{\omega}\mu_{0})^{2}}{16\pi^{2}}\,\frac{2a}{z_{1}z_{2}}\;e^{+j\beta\,\Phi_{12}}\;\mathrm{sinc}\left(\beta\frac{(R_{2}^{+}-R_{2}^{-})-(R_{1}^{+}-R_{1}^{-})}{2}\right),$ (31)
$\Phi_{12}$ $=\frac{R_{2}^{+}+R_{2}^{-}}{2}-\frac{R_{1}^{+}+R_{1}^{-}}{2},$ (32)
where $\mathrm{sinc}(x)\equiv\sin x\,/\,x$.
The magnitude of the spectral degree of coherence
$\big|\mu(\mathbf{r}_{1},\mathbf{r}_{2})\big|\;\approx\;\mathrm{sinc}\left(\beta\frac{(R_{2}^{+}-R_{2}^{-})-(R_{1}^{+}-R_{1}^{-})}{2}\right),$ (33)
depends only on the difference in optical path lengths from the two aperture edges to the observation points. [^18] The connection with the Geometrical Theory of Diffraction follows from the close analogy between coherence (CSD) integrals and classical diffraction integrals.
Consequently $|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|$—and hence the mutual information—is constant along the entire hyperbola (within the given approximation) $R_{2}^{+}-R_{2}^{-}=\text{const}$ passing through $\mathbf{r}_{1}$ . This means that for a fixed point $\mathbf{r}_{1}$, the field at any point $\mathbf{r}_{2}$ on the hyperbola passing through it is highly predictable. These curves, along which the correlation is high, are referred to as "Spatial Information Flows" (SIFs), a concept introduced in [40, 41].
SIFs, therefore, coincide with the curves of constant $\xi$ defined in [18]. Additionally, the mutual information is almost null between any two points belonging to different hyperbolas in the set $R_{2}^{+}-R_{2}^{-}=n\lambda$. This subset of SIFs will be called $Q_{SIF}$ and coincides  [^19] As noted in the footnote of Section 3.2, on observation surfaces the sampling approach can yield a slightly larger number of samples compared to the $N_{DoF}\,\,$, because achieving a prescribed approximation accuracy typically requires a modest enlargement of the usable bandwidth with respect to the effective bandwidth introduced in [17]. However, if the sampling is chosen consistently with the field’s effective bandwidth value, then $Q_{\xi}$ and $Q_{\mathrm{SIF}}$ coincide in 2D geometry. Loosely speaking, the effective bandwidth marks the boundary between packing and covering descriptions of the received–field set [5]. with the subset $Q_{\xi}$. The equivalence of these two sets, one defined in approximation theory and the other in EIT, is a consequence of the fact that they both represent the same physical limitations in different domains of application.
From a ’spatial information’ perspective, the linear source effectively perceives its surroundings as a hyperbolically warped space  [^20] The warping is given by the conformal transformation $w(z)=\operatorname{arccosh}\left(-{i\zeta}/a\right)$, with the complex variable $\zeta=y+iz$. The approach in [42] implicitly identifies the warping transformation for curved observation domains by working directly on the radiation operator. The identification of warp transformations for surface observation domains remains a challenging and active area of research [43]., as illustrated in Fig. 3. The transformation linearizes the hyperbolas (red curves) and ellipses (blue curves) of Fig. 2. The transformation linearizes the hyperbolas (red curves) and ellipses (blue curves) of Fig. 2. This new geometry provides a clear visualization of how the Degrees of Freedom are distributed across the different receiving domains, effectively modifying the observation-domain geometry perceived by the antenna.
The above discussion is based on an approximated expression of the correlation. To identify the actual zone around a point $\mathbf{r}_{1}$ where mutual information keeps high, we will numerically estimate of the CSD and the mutual information.
Figure 5: Spectral degree of coherence amplitude, $|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|$ (left), and mutual information, $I_{MI}(\mathbf{r}_{1},\mathbf{r}_{2})$ (right), for a linear source of length $\ell=8\lambda$. The point $\mathbf{r}_{1}$, indicated as $\bar{\mathbf{r}}$ in the main text, is marked with a white circle, while $\mathbf{r}_{2}$ ranges across the entire domain of the figure. The subset $Q_{\xi}$ of hyperbolas of constant $\xi$ is drawn as dashed curves.
Figure 6: Spectral degree of coherence amplitude, $|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|$ (left), and mutual information, $I_{MI}(\mathbf{r}_{1},\mathbf{r}_{2})$ (right) for a linear source of length $\ell=8\lambda$. The point $\mathbf{r}_{1}$ is marked with a white circle, while $\mathbf{r}_{2}$ ranges across the entire domain of the figure. The subset $Q_{\xi}$ of hyperbolas of constant $\xi$ is drawn as dashed curves.
Before presenting our results, we first examine the Wigner Distribution Function (WDF). Figure 4 shows the real part of the WDF along the hyperbola that passes through the point $\mathbf{r}_{1}=\bar{\mathbf{r}}$ where $\bar{\mathbf{r}}\simeq(20\lambda,16\lambda)$. The plot is a function of the normalized curvilinear abscissa $s/\lambda$ and of the component of the wavevector tangent to the hyperbola, $k_{s}$. The value of $s/\lambda$ in $\bar{\mathbf{r}}$ is indicated by a horizontal dashed line.
The figure shows that the angular spectrum in $\mathbf{r}_{1}=\bar{\mathbf{r}}$ is strongly concentrated towards $k_{s}=\beta$. This indicates that the hyperbola is aligned with the optical ray passing through $\bar{\mathbf{r}}$ . For all points to the right of the dashed line, the hyperbola tends toward a straight line, and the $\xi$ curve coincides with the optical ray passing through $\bar{\mathbf{r}}$. In this region, the cross-spectral density (CSD) function locally approximates that of a coherent plane wave, and the hyperbola also coincides with the SIF passing through $\bar{\mathbf{r}}$.
Figures 5 display $|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|$ (left), and $I_{MI}(\mathbf{r}_{1},\mathbf{r}_{2})$ (right), for the same scenario as in Figure 2. The plots also show the $Q_{\xi}$ curves as white dotted lines. The fixed point $\mathbf{r}_{1}=\bar{\mathbf{r}}$ is marked with a white circle, while $\mathbf{r}_{2}$ ranges across the entire domain of the figure. The figure confirms the emergence of a high-mutual-information path, i.e., a SIF, which is centered on the hyperbola.
To complete the discussion, we repeat the same analysis, this time moving $\mathbf{r}_{1}$ closer to the source. The new location of $\mathbf{r}_{1}$ is indicated by the dotted vertical line in Fig. 4. The real part of the WDF shows a broader angular power spectrum, which is typical of the Fresnel zone.
The point $\mathbf{r}_{1}$ is marked with a white circle in Figs. 6. These figures show that both $|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|$ and the mutual information still follow the hyperbolas (i.e., the constant $\xi$ curves) in the Fresnel region, although the region of high mutual information significantly shrinks.
While a full analysis of information flow in the Fresnel zone requires further investigation, these results suggest that the SIFs still coincide with the constant $\xi$ curves. This observation further corroborates the effectiveness of deterministic models in EIT.
5 Conclusions
This paper investigates the relationship between deterministic and stochastic source models in Electromagnetic Information Theory (EIT), using tools from optical coherence theory.
Our findings demonstrate a fundamental connection between these two approaches. We show that deterministic models and stochastic models with spatially homogeneous uncorrelated sources yield identical eigenvalues and basis functions for field representation. This equivalence explains the effectiveness of deterministic approaches in EIT and corroborates the use of classical electromagnetic methods within the EIT framework.
In essence, this result is both simple and profound: it shows that deterministic models are sufficient to fully capture the physical limitations underlying EIT.
Recalling the metaphor from [39], where electromagnetics is the horse and information theory is the horseman, the horse ultimately determines the limits of a communication system, regardless of the riders. Getting the best out of the horse to truly revolutionize communication systems is the challenge of Electromagnetic Signal and Information Theory (ESIT) [44].
As a final remark, in this article we have briefly noted that the $N_{DoF}\,\,$in communication systems can be estimated under certain conditions using principles from radiometry. This observation indicates that the $N_{DoF}\,\,$is a metric intrinsically linked to the physics of energy transfer. Understanding the relationships between information and energy in the context of the $N_{DoF}\,\,$would represent another small piece in our understanding of how the interplay between information and energy shapes our universe.
Appendix A Appendix: Equivalence of Coherent Modes and Singular Functions for Spatially Incoherent Sources
Notation.
In the following, the symbol $\dagger$ denotes the conjugate transpose (Hermitian transpose) for finite-dimensional quantities such as vectors and dyadics, and the Hilbert-space adjoint for operators with respect to the $L^{2}$ inner product.
Hypothesis
Let $D\subset\mathbb{R}^{3}$ be a bounded, measurable source domain and $S_{\mathrm{obs}}\subset\mathbb{R}^{3}$ a measurable observation domain. Consider a linear, time-invariant (LTI) medium characterized by a dyadic Green’s function $\bar{\bar{\mathbf{G}}}(\mathbf{r},\mathbf{r}^{\prime})\in L^{2}(S_{\mathrm{obs}}\times D;\mathbb{C}^{3\times 3})$.
  1. 1.
    The Radiation Operator: Define the Hilbert–Schmidt (and therefore compact) linear integral operator $\mathcal{A}:L^{2}(D;\mathbb{C}^{3})\to L^{2}(S_{\mathrm{obs}};\mathbb{C}^{3})$ as:
    $(\mathcal{A}\mathbf{J})(\mathbf{r})=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r},\mathbf{r}^{\prime})\mathbf{J}(\mathbf{r}^{\prime})\,d\mathbf{r}^{\prime},\qquad\mathbf{r}\in S_{\mathrm{obs}}.$
    Let $\{\mathbf{u}_{n}\}_{n\geq 1}$ be the set of orthonormal left singular functions and $\{\sigma_{n}\geq 0\}_{n\geq 1}$ the corresponding singular values of $\mathcal{A}$, defined by the eigenvalue problem:
    $\mathcal{A}\mathcal{A}^{\dagger}\,\mathbf{u}_{n}=\sigma_{n}^{2}\,\mathbf{u}_{n}.$
  2. 2.
    The Stochastic Source: Let the source $\mathbf{J}(\mathbf{r}^{\prime})$ be a zero-mean, spatially delta-correlated random vector field within $D$. Its cross-spectral density (CSD) dyadic is
    $\bar{\bar{\mathbf{W}}}_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=\mathbb{E}\big[\mathbf{J}(\mathbf{r}^{\prime}_{1})\mathbf{J}(\mathbf{r}^{\prime}_{2})^{\dagger}\big]=\bar{\bar{\mathbf{I}}}\,\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2}),$
    where $\bar{\bar{\mathbf{I}}}$ is the $3\times 3$ identity dyadic and $\mathbb{E}[\cdot]$ denotes statistical expectation.
  3. 3.
    The Radiated Field: The random electric field $\mathbf{E}(\mathbf{r})$ on $S_{\mathrm{obs}}$ results from the linear propagation $\mathbf{E}=\mathcal{A}\mathbf{J}$. Its statistical properties are described by the field CSD kernel $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})=\mathbb{E}\big[\mathbf{E}(\mathbf{r}_{1})\mathbf{E}(\mathbf{r}_{2})^{\dagger}\big]$. This kernel defines a positive semidefinite integral operator $\mathcal{W}$, whose orthonormal eigenfunctions $\{\bm{\psi}_{n}\}$ (the coherent modes) and non-negative eigenvalues $\{\lambda_{n}\}$ are the solutions to the eigenvalue problem:
    $(\mathcal{W}\bm{\psi}_{n})(\mathbf{r}_{1})=\int_{S_{\mathrm{obs}}}\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})\,\bm{\psi}_{n}(\mathbf{r}_{2})\,d\mathbf{r}_{2}=\lambda_{n}\,\bm{\psi}_{n}(\mathbf{r}_{1}).$
Thesis
The set of coherent modes $\{\bm{\psi}_{n}\}$ of the radiated random field is identical (up to a global phase factor) to the set of left singular functions $\{\mathbf{u}_{n}\}$ of the radiation operator $\mathcal{A}$. Their corresponding eigenvalues are equal to the squared singular values:
$\bm{\psi}_{n}(\mathbf{r})=e^{i\phi_{n}}\,\mathbf{u}_{n}(\mathbf{r}),\qquad\lambda_{n}=\sigma_{n}^{2},\qquad\forall\,n\geq 1,$
where $e^{i\phi_{n}}$ represents an arbitrary global phase factor for each mode.
Demonstration
The proof shows that the CSD kernel $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})$ is identical to the kernel of the operator $\mathcal{A}\mathcal{A}^{\dagger}$.
  1. 1.
    Compute the field CSD kernel $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})$ by propagating the source CSD:
    $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})$ $=\mathbb{E}\big[(\mathcal{A}\mathbf{J})(\mathbf{r}_{1})(\mathcal{A}\mathbf{J})(\mathbf{r}_{2})^{\dagger}\big]$
    $=\mathbb{E}\!\left[\left(\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})\mathbf{J}(\mathbf{r}^{\prime}_{1})\,d\mathbf{r}^{\prime}_{1}\right)\left(\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})\mathbf{J}(\mathbf{r}^{\prime}_{2})\,d\mathbf{r}^{\prime}_{2}\right)^{\dagger}\right]$
    $=\iint_{D\times D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})\,\mathbb{E}\!\big[\mathbf{J}(\mathbf{r}^{\prime}_{1})\mathbf{J}(\mathbf{r}^{\prime}_{2})^{\dagger}\big]\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})^{\dagger}\,d\mathbf{r}^{\prime}_{1}\,d\mathbf{r}^{\prime}_{2}$
    $=\iint_{D\times D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})\,\big(\bar{\bar{\mathbf{I}}}\,\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2})\big)\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})^{\dagger}\,d\mathbf{r}^{\prime}_{1}\,d\mathbf{r}^{\prime}_{2}$
    $=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\,d\mathbf{r}^{\prime}.$
  2. 2.
    Find the integral kernel of the operator $\mathcal{A}\mathcal{A}^{\dagger}$. For any $\mathbf{f}\in L^{2}(S_{\mathrm{obs}};\mathbb{C}^{3})$:
    $(\mathcal{A}\mathcal{A}^{\dagger}\mathbf{f})(\mathbf{r}_{1})$ $=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,(\mathcal{A}^{\dagger}\mathbf{f})(\mathbf{r}^{\prime})\,d\mathbf{r}^{\prime}$
    $=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\!\left(\int_{S_{\mathrm{obs}}}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\mathbf{f}(\mathbf{r}_{2})\,d\mathbf{r}_{2}\right)d\mathbf{r}^{\prime}$
    $=\int_{S_{\mathrm{obs}}}\left(\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\,d\mathbf{r}^{\prime}\right)\mathbf{f}(\mathbf{r}_{2})\,d\mathbf{r}_{2}.$
    By inspection, the kernel of $\mathcal{A}\mathcal{A}^{\dagger}$, let’s call it $\bar{\bar{\mathbf{K}}}_{\mathbf{u}}(\mathbf{r}_{1},\mathbf{r}_{2})$, is
    $\bar{\bar{\mathbf{K}}}_{\mathbf{u}}(\mathbf{r}_{1},\mathbf{r}_{2})=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\,d\mathbf{r}^{\prime}.$
  3. 3.
    Conclusion: Comparing the results of steps 1 and 2 shows the kernels are identical, $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})\equiv\bar{\bar{\mathbf{K}}}_{\mathbf{u}}(\mathbf{r}_{1},\mathbf{r}_{2})$, which implies the operators themselves are identical:
    $\mathcal{W}\equiv\mathcal{A}\mathcal{A}^{\dagger}.$
    Since the operators are the same, their respective eigenvalue problems, $\mathcal{W}\bm{\psi}_{n}=\lambda_{n}\bm{\psi}_{n}$ and $(\mathcal{A}\mathcal{A}^{\dagger})\mathbf{u}_{n}=\sigma_{n}^{2}\mathbf{u}_{n}$, must be equivalent. It follows that their eigenfunctions, $\bm{\psi}_{n}$ and $\mathbf{u}_{n}$, form the same orthonormal basis (up to arbitrary phase factors), and their eigenvalues are equal, $\lambda_{n}=\sigma_{n}^{2}$. This completes the proof.
Acknowledgment
The author is indebted to Professor Ovidio Mario Bucci for their insightful discussions and valuable suggestions, which enriched the quality and rigor of this work.